Question 6 Find the volume formed by rotating the region enclosed by: x = 4.4y and y = x with y > 0 about the y-axis
Answer #1

solution: - Given that a= 4.49 y=x.eso WWNA -7x yzo. n24049 y = 4.44 y (44–4.4)50 you y = 4.4214 - yz 22 х» 4. 49 249 *Vaz =